### v1

# Can Neural Decompilation Assist Vulnerability Prediction on Binary Code?

Abstract.
Vulnerability prediction is valuable in identifying security issues more efficiently, even though it requires the source code of the target software system, which is a restrictive hypothesis. This paper presents an experimental study to predict vulnerabilities in binary code without source code or complex representations of the binary, leveraging the pivotal idea of decompiling the binary file through neural decompilation and predicting vulnerabilities through deep learning on the decompiled source code. The results outperform the state-of-the-art in both neural decompilation and vulnerability prediction, showing that it is possible to identify vulnerable programs with this approach concerning bi-class (vulnerable/non-vulnerable) and multi-class (type of vulnerability) analysis.
Vulnerability Prediction, Binary Analysis, Neural Decompilation, Security, Deep Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineeringccs: Security and privacy Vulnerability management
1. Introduction
Vulnerability prediction is an approach to analyze and predict which files or functions of a software system are most likely vulnerable. This approach enables focusing effort and resources (e.g., man-hours) on the vulnerable parts. However, vulnerability prediction has the drawback of being a source code-based approach, which is a restrictive hypothesis: source code is rarely available for several classes of software systems, such as proprietary software, legacy systems, or firmware monitoring specific hardware. Limited access to source code hinders vulnerability prediction, hence the need for new solutions to identify vulnerabilities when source code is unavailable (yan2021han, ; yamaguchi2012generalized, ; xu2018vulnerability, ).
Table 1. Intermediate steps from binary file to vulnerability prediction.
Binary ASM x86 Decompiled C++ Vulnerability
0101 1001 0111 ... <funct0>: endbr 64 push rbp ... void funct0(){ int data; data=0; ... • The function is vulnerable.
• The vulnerability is linked to CWE 190.
Neural decompilation is an emerging method that leverages modern deep learning techniques, such as transformers and self-attention, to translate binary code back into source code. One of the key benefits of this approach is that the control flow structure of the decompiled program resembles the original source code, in a closer way than disassemblers and traditional decompiler tools, such as Ghidra (Ghidra, ). It is a data-driven approach, where the neural network learns the typical patterns of source code structures. Gaining access to the source code offers several advantages, such as enabling vulnerability detection, malware recognition, re-engineering of legacy systems, and facilitating static analysis. Moreover, binary decompilation enables vulnerability prediction in “black-box” vulnerability assessments (e.g., proprietary software), where the assessor attacks the system from the same perspective of the attacker, without source code.
Figure 1. Overview of vulnerability prediction via neural decompilation.
In recent years, several solutions have emerged to predict vulnerabilities through binary code analysis, which can analyze a binary file even if its source code is unavailable (dahl2020stack, ; eschweiler2016discovre, ; grieco2016toward, ; pewny2015cross, ; schaad2023deep, ; zheng2019recurrent, ). These solutions rely on assembly code obtained from disassemblers, or on intermediate code representations such as the one from the LLVM toolchain (hasabnis2016lifting, ). However, these representations do not fully capture the semantic richness provided by the source code. The difficulties lie in several aspects: i) lack of structural information about the layout of data and code, which can be interleaved; ii) complexity and heterogeneity of the target architectures (which may have highly variable instruction sets and calling conventions); iii) optimizations performed by the compiler, which can significantly modify the program structure, (e.g., loop optimizations); and iv) lack of function and variable identifiers, which are typically stripped away from commercial software binaries (arakelyan2021bin2vec, ). Moreover, binary-oriented solutions, such as traditional decompilers, present intrinsic limitations due to the semantic gap between source and binary code (e.g., loss of the control flow structure during compilation).
In this work, we prove the potential of identifying software vulnerabilities in binary files via neural decompilation (Figure 1), without needing to access the source code of the target software. Table 1 shows the output of each step of the proposed approach. The results of this study show that:
  1. (1)
    Neural decompilation achieves impressive performance in the decompilation of C/C++ code, reaching up to an Edit Distance of 59% with the fairseq model.
  2. (2)
    Vulnerability prediction yields highly accurate results:
    • the CodeGPT model successfully determines whether the code is vulnerable (bi-class analysis, 95% F1-Score).
    • the CodeBERT model accurately predicts (83% F1-Score) the type of vulnerability among 20 different vulnerabilities (multi-class analysis), mapped onto the Common Weaknesses Enumeration (CWE) (CWE, ) defined by MITRE (Mitre, ).
  3. (3)
    Decompiled source code leads to better vulnerability prediction accuracy compared to disassembled code (schaad2023deep, ), showing a 7% improvement in the bi-class analysis and a 6% improvement in the multi-class analysis.
The remainder of the paper is structured as follows. In Section 2, we elaborate on background concepts about the target problem. In Section 3, we present the experimental approach for vulnerability prediction via neural decompilation. In Section 4, we illustrate the experimental analysis for the decompilation and vulnerability prediction tasks. In Section 5, we elaborate on the experimental evaluation results. In Section 6, we examine related work. In Section 7, we discuss the threats to validity. Section 8 concludes the paper.
2. Background
The main complexities of this work lie in the decompilation process, i.e., the transformation of executable code into source code, which is not completely reversible. Several decompilers are available in this field (e.g., Ghidra (Ghidra, ), RetDec (RetDec, )), achieving good levels of accuracy, however, they come with high costs in terms of knowledge of the specific language, and generation of control flow graphs. Traditional decompilation methods typically leverage an intermediate language (IL), independent from the architecture used to trace the source code, which increases the cost of developing such a system even more. In addition, the resulting decompiled files are not highly readable, as they closely reflect the structure of the assembly code.
2.1. Semantic Gap
To understand the issues involved in decompilation, we recap the steps of the compilation process, from a program written in a high-level language to the corresponding executable file. First, the program is compiled into an assembly file. Next, the assembly is transformed into a binary object file, i.e., machine code not yet executable because of unresolved function calls reported in the relocation information. Therefore, the linker must transform the object file into an executable file by linking standard library functions and other programs.
In contrast, transforming executable code into source code requires different methodologies and tools. The difficulties lie in the operations of disassembly and decompilation. Disassembly consists of lifting binary code to assembly code. There are several open-source tools available for this task. On the other hand, there is not a predefined reverse procedure for decompilation: once a program is compiled into low-level code most of the information is lost.
This issue stems from the semantic gap between high- and low-level code: machine code does not provide function-level abstraction since function calls are implemented through call labels and jumps that allow the execution of loops and conditional constructs. This asymmetry causes the loss of the control flow of the original program, which is quite challenging to re-construct from the binary.
Another problem is that compilers may rearrange code differently and transform it according to optimization rules. For example, a compiler may implement multiplication/division through a series of sums/subtractions. Also, an assembled file typically contains operating system directives that do not affect the particular program. Unaware of such complications, a decompiler would only translate what is reported in the assembly code.
2.2. Dataset
One of the complexities in dealing with the translation and vulnerability prediction tasks lies in finding an appropriate dataset. The dataset should meet the following requirements:
  • Compilable: the programs under analysis in the dataset must be compilable to generate executables for the binary analysis.
  • Standardized classification: the programs must be labeled according to the vulnerability they represent following an appropriate categorization system, i.e., a standardized vulnerability taxonomy.
In addition, we will need to pre-process the collection of programs before feeding the decompilation model. The output of such a model will be the input to the vulnerability prediction stage.
Table 2. C/C++ datasets for vulnerability prediction.
Dataset # Samples Compilable Standardized Taxonomy
ReVeal (arewethereyet, ) $18,169$ No N/A
Devign (zhou2019devign, ) $26,037$ No N/A
Juliet (Juliet, ) $128,198$ Yes Yes
BigVul (bigvul, ) $264,919$ No Only for 8,783 samples
DiverseVul (chen2023diversevul, ) 330,492 No Only for 16,109 samples
Table 2 presents an overview of the C/C++ datasets for vulnerability prediction available in the literature. It is possible to notice that datasets such as BigVul (bigvul, ) or DiverseVul (chen2023diversevul, ) provide a significant number of samples. However, they do not provide compilable functions, hence violating the first dataset requirements, i.e., compilability. On the other hand, they map the vulnerabilities on the CWEs taxonomy for small fractions of the samples. ReVeal (arewethereyet, ) and Devign (zhou2019devign, ) do not fulfill both requirements: their C/C++ code is not compilable and their vulnerabilities are not mapped on any standardized taxonomy. For these reasons, the only suitable dataset for our study is Juliet C/C++ 1.3 (Juliet, ), a synthetic collection of C/C++ programs from the Software Assurance Reference Dataset (SARD) (SARD, ), developed by NIST (NIST, ). The Juliet test suite was first released in 2010, gaining increasing popularity for benchmarking static analysis tools and vulnerability prediction tasks (zheng2019recurrent, ; schaad2023deep, ; li2021vuldeelocator, ; li2018vuldeepecker, ; jain2023code, ; gao2018comprehensive, ) since it includes both vulnerable and non-vulnerable programs. The structure of the dataset provides:
  • $64,099$ compilable test cases for 118 CWEs. For each CWE, a test case is composed of a program vulnerable to the target CWE and a non-vulnerable one.
  • A variable number of test cases for each CWE, ranging from a few units to thousands of examples.
  • A single test case may consist of a single C/C++ file or have a modular structure consisting of multiple files, with header files containing the functions’ definitions.
  • When compiled, each test case generates two executables, a bad and a good version, respectively, a version that does not handle the vulnerability and one that implements preventive measures.
3. Methodology
Figure 2. Experimental approach for our study.
Before presenting the experimental approach, we introduce the following assumptions:
  • To avoid translating binary code to high-level code directly, we disassembled the binary into assembly code. Consequently, the input for the decompilation model will be assembly code, further pre-processed in the next steps.
  • Assembly code is defined by the architecture used to generate the executables. In our experimental evaluation, the assembly language is for the x86 architecture.
  • The high-level target language of the translation is C/C++. These languages are still widely used to develop network applications, Internet of Things (IoT) systems, and systems with safety requirements.
  • The neural decompilation model must learn to translate code at the function level. Through pre-processing, we strip the programs of unnecessary information associated with libraries, namespaces, external modules, or main functions that only invoke other functions.
Our experimental approach revolves around five phases: Binary Generation, Disassembly, Pre-processing, Neural Decompilation, and Vulnerability Prediction, as in Figure 2.
3.1. Binary Generation
This stage consists of compiling the programs under analysis to generate two executables for each of them, a bad and a good version (JulietGithub, ). Respectively, bad executables are vulnerable to the target CWE, while good executables implement countermeasures to prevent it.
As introduced in Section 2.2, Juliet contains $64,099$ samples that cover 118 CWEs, however, a subset of such programs is not compilable due to issues with external libraries. It is worth noting that the programs were compiled without specific optimization options (JulietGithub, ), as performed by other works in the field (schaad2023deep, ). Consequently, the resulting test cases ($41,760$) cover 89 CWEs (75% of the total number) that span several categories of vulnerabilities, e.g., memory and resource management, injection flaws. Table 3 reports some statistics about the compiled dataset.
3.2. Disassembly
This stage lifts the binary code to assembly code. This transformation depends on the target architecture for executables. Several tools are available for this task: we chose objdump (objdump, ), a popular and easy-to-use tool. The format of the executables is elf64-x86-64.
3.3. Pre-processing
Assembly code. The machine code files are not yet ready for training since they are extremely long and contain library functions irrelevant to our purpose. An assembly file consists of several sections, with the code section being .text. We removed all sections except .text since we are only interested in the algorithmic logic.
At this stage, it is still necessary to remove library functions defined within the same section, operating system directives, stack initialization, and the main function, which is constant in every program, making its contribution to learning negligible. What is left are exclusively the (non-)vulnerable functions of interest. We aim to let the model learn algorithm recognition and structural patterns related to functional programming logic. We also replaced function names with standard names for each test case. This operation allows the model to focus on the function content rather than the function name.
The main benefit of this stage concerns the length of the assembly code for each test case, which is drastically reduced by more than 90% (on average). Then, the assembly code is ready for the translation process.
High-level code. High-level language programs require both similar and additional pre-processing operations to those performed on the assembly code. First, printed strings can be removed from the target programs since there is no solution for the model to predict them. Since we know that each program generates its own bad and good versions, one of the fundamental operations to perform is splitting them. For this purpose, we leveraged pre-compiler directives.
The next pre-processing step is comment removal. Similarly to low-level code pre-processing, we removed the main function and the inclusions of libraries and namespaces. Again, we replaced function names with standard names (e.g., funct0) to let the model focus exclusively on the algorithm of the actual functions.
Tokenization. After the pre-processing phase, both the high- and low-level code are represented as unstructured text, hence the need to tokenize it. The first stage of this process is a pre-tokenization to split the words of the training data according to the blanks. We adopted the Byte Pair Encoding (BPE) technique, typically preferred for tasks such as NMT (hosseini2022beyond, ). Then, we performed toss reduction, i.e., removal of programs that exceed a specific length in terms of the number of tokens: we cut off those programs whose size is above the 95th percentile ($1,479$ tokens for the source language, $338$ for the target) and below the 5th percentile ($20$ for the source language, $9$ for the target). As a result of this operation, the number of elements decreased from $83,520$ to $75,149$.
Table 3. Dataset pre-processing
Statistic Compiled Dataset Pre-processed Dataset
C/C++ ASM C/C++ ASM
Function Count $83,520$ $83,520$ $75,149$ $75,149$
CWE Count $89$ $89$ $89$ $89$
Avg. Function Length $612$ $1,821$ $87$ $158$
Max Function Length $1,148$ $17,680$ $338$ $1,479$
3.4. Neural Decompilation
Neural decompilation is an emerging approach to translating binary code into source code. The main advantages of this approach are the following:
  • The decompiled program control-flow structure traces the structure of the original high-level program, making the results human-readable. This happens because the transformer learns the typical structure of source programs.
  • Neural decompilation is a data-driven approach, hence it can be trained on new target languages or architectures with new samples set without depending on specific and hard-to-use toolchains.
Neural decompilation leverages Neural Machine Translation (NMT), a machine learning approach to translating a given sequence of elements into another sequence through the use of Large Language Models (LLMs), deep learning algorithms able to perform a set of language processing tasks such as translation, classification or generation. In light of the complexity of the tasks, LLMs generally require large accurate datasets to be trained. The core of the LLMs is the Transformer neural network, a deep learning architecture consisting of an encoder-decoder structure based on the multi-headed attention mechanism (vaswani2017attention, ), often used for translation from one language to another. In our case, we leverage NMT to translate machine code to high-level code. The main difference between the two uses is the number of tokens between the source and the translated text. There is a wide gap in the number of tokens between the machine and high-level code. Another difference is that, although natural language follows grammatical rules, a programming language is more structured by nature. The fine-tuned models for the translation task are the following:
  • CodeBERT (feng2020codebert, ): is a large multi-layer bidirectional Transformer with an encoder-only architecture pre-trained on millions of lines of code across six different programming languages created by Microsoft. For the translation task, we combined it with a decoder to have an encoder-decoder architecture suitable for code generation (wang2021codet5, ).
  • CodeT5+ (wang2023codet5, ): is a large multi-layer bidirectional Transformer with an encoder-decoder architecture, developed by Salesforce and pre-trained on diverse tasks including causal language modeling, contrastive learning, and text-code matching. We utilize the variant with model size $220M$, trained from scratch following T5’s architecture (wang2021codet5, ) and initialize it with a checkpoint further pre-trained on Python.
  • fairseq (fairseq, ): is a sequence modeling toolkit written in PyTorch that allows researchers and developers to train custom models for text translation and generation, created by Meta. This model is not pre-trained and supports encoder-decoder architectures for a wide variety of tasks. We selected this model on the basis of previous work (hosseini2022beyond, ).
3.5. Vulnerability Prediction
This stage represents the core of the vulnerability prediction task. The goal is to identify whether a target program is vulnerable, along with the type of vulnerability affecting it. We addressed these problems as classification problems, following the direction of several studies about vulnerability prediction (zheng2019recurrent, ; schaad2023deep, ; pewny2015cross, ). Specifically, we faced the following classification tasks:
  • Binary classification: deals with finding whether a target C/C++ function is vulnerable (bad) or not (good), regardless of the specific vulnerability.
  • Multi-class classification: aims to identify which vulnerability affects the target program. Vulnerability types are the classes to be predicted by the classification models. Such vulnerabilities are mapped to the CWE (CWE, ), a standard taxonomy defined by MITRE (Mitre, ).
Therefore, we train several models on a training set consisting of original C/C++ programs, i.e., non-decompiled, and then test them on a test set consisting of neurally decompiled C/C++ programs. We make sure that overfitting does not occur since none of the samples in the original C/C++ training set are contained in the decompiled test set. Looking at previous studies on this topic (chen2023diversevul, ; schaad2023deep, ; zheng2019recurrent, ), we selected models that can handle the complexity of vulnerability prediction, including:
  • CodeBERT (feng2020codebert, ): unlike its use for the neural decompilation task, we used it as an encoder-only architecture for vulnerability prediction.
  • CodeT5+ (wang2023codet5, ): described in Section 3.4.
  • CodeGPT (huggingface2020codegpt, ): a Transformer based on the GPT-2 model family. CodeGPT provides a decoder-only architectural configuration. It was primarily designed for code generation and optimized for Python, but can also be adapted to generic code classification tasks.
  • Long Short-Term Memory (LSTM): a recurrent neural network with retroactive connections that provides short-term memory. This makes it fit for processing long input data sequences, as in the case of code classification.
  • Simple Recurrent Neural Networks (SRNN): it consists of a neural network with interconnections between units through cycles. This enables determining a temporal dynamicity, dependent on the information received at the previous times.
  • Gated Recurrent Unit (GRU): consists of a recurrent neural network that introduces a gating mechanism, i.e., a mechanism to decide which information is forwarded to the output.
Table 4. Automatic evaluation of the neural decompilation task. Best performance is blue.
Approach Model # functions Max target length BLEU-4 ED METEOR ROUGE-L
*Ghidra (Ghidra, ) IR 1,000 338 22% 23% 29% 22%
°Katz et al. (katz2018using, ) RNN 700,000 88 - 30% - -
°Hosseini et al. (hosseini2022beyond, ) fairseq 2,000,000 271 - 54% - -
Proposed approach CodeBERT 75,000 338 31% 39% 43% 56%
CodeT5+ 75,000 338 48% 49% 55% 70%
fairseq 75,000 338 58% 59% 64% 77%
*: results obtained on a random test set of Juliet dataset.
°: trained with different data than Juliet.
4. Experimental Analysis
4.1. Experimental Setup
The analysis was performed on a virtual instance of Google Cloud running on Debian GNU/Linux OS version 10, equipped with an Intel Haswell CPU, 16 GB RAM, a 250 GB disk, and a Tesla V100-SXM2-16GB GPU.
Neural Decompilation. The translation task was performed using fairseq, CodeBERT, and CodeT5+ (base: 220M) models. CodeBERT is trained by implementing an encoder-decoder framework where the encoder is initialized with the pre-trained CodeBERT weights, while the decoder is a transformer decoder comprising 6 stacked layers. The encoder leverages the RoBERTa architecture (liu2019roberta, ), with 12 attention heads, 768 hidden layers, 12 encoder layers, and 514 for the size of position embeddings.
CodeT5+ has an encoder-decoder architecture with 12 decoder layers, each with 12 attention heads and 768 hidden layers, and 512 for the size of position embeddings. We set learning rates of 5e-5, batch size of 8, beam size of 10, evaluation steps of 314, and training steps of 2005 for both CodeBERT and CodeT5+ models. Source and target length were set to 512, the maximum length of the input sequences in both CodeBERT and CodeT5+ models. We truncated all the assembly programs longer than the maximum supported length.
We derived basic fairseq hyperparameters configuration for model training from the work of Hosseini et al. (hosseini2022beyond, ). We chose a different optimizer, namely the nag optimizer. We performed 10-fold cross-validation to evaluate the decompilation models using the whole dataset and have a more robust and reliable evaluation. Specifically, we trained and evaluated the models ten times, each using one of the ten folds as the test set and the remaining nine as the training set. At the end of the ten iterations, performance is averaged to obtain a more reliable estimate of the model’s predictive capabilities. Vulnerability Prediction. For the neural network-based models, namely LSTM, SRNN, and GRU, we used 4 bidirectional layers with a number of units equal to the maximum sequence length set during pre-processing and one dense layer with a number of layers equal to the number of classes for all three networks. For training, we employed the Categorical Cross-entropy metric and the Adam optimizer, with an initial learning rate of $41e-$ 4. We configured $100$ epochs, adding callbacks to improve and speed up learning. We set the patience value for Early Stopping to 5 epochs, hence the training was stopped if the loss did not grow within 5 epochs. Each training lasted between $30$ and $40$ epochs. We chose the Word2Vec Gensim library for pre-processing, i.e., from tokens to embeddings. We implemented input padding to prevent the model from basing its choices on the programs length.
On the other hand, for the transformer-based models, we set the batch size per device to 4 for both training and evaluation. The learning rate was set to $52e-$ 5, a low value to ensure gradual updates, employing a linear learning rate scheduler to progressively reduce the learning rate throughout training. The weight decay was set to 0.01 to regularize the model and prevent overfitting. The Adam optimizer was used with an epsilon of $81e-$ 8 to ensure numerical stability. We applied gradient accumulation in two steps, which would mean the gradients were accumulated in two iterations before updating the model. The models were trained for 3 epochs. We utilized the tokenizer models provided by the LLMs themselves for proper input processing: T5Tokenizer for T5, GPT2Tokenizer for CodeGPT, and RobertaTokenizer for CodeBERT.
4.2. Evaluation - Neural Decompilation
We evaluated the translation quality using output similarity metrics including BLEU, METEOR, ROUGE, and Edit Distance (ED), illustrated in Table 3.5. Each metric expresses different characteristics of translation quality: BLEU allows capturing the precision of n-grams, giving importance to sequence length; ROUGE is useful in capturing code structure and overlapping semantic units, such as word sequences and key phrases; METEOR takes into account precision, completeness, and word order. However, metrics such as BLEU, originally developed for natural text translation, may not be particularly appropriate for evaluating translation between code sequences (tran2019does, ; liguori2023evaluates, ). ED is suitable for capturing small variations in the code and is representative of the number of changes to perform on the decompiled output. Normalized ED lies between 0 and 1 and for each $i_{th}$ sequence pair $(hyp_{i}-ref_{i})$ is calculated as: Normalized ED_i = 1 - ED(hypi,refi)max(len(refi),len(hypi)) Hence, a high Normalized ED value corresponds to an optimal translation. In the following, we refer to this metric as ED to maintain coherence with other works in this field.
Table 3.5 compares different decompilation solutions based on their approach, number of functions used for training and testing, maximum target length, and output similarity metrics. The results of neural decompilation are outstanding. Specifically, fairseq achieves a 59% ED score using about $75,000$ C/C++ functions for both training and testing with a maximum target length of 338 tokens. These results are close to those of other studies on this topic: Hosseini et al. (hosseini2022beyond, ) reached an 54% ED using about $2,000,000$ C/C++ functions with a maximum length of $271$, while Katz et al. (katz2018using, ) achieves an ED of 30% using about $700,000$ C-code snippets with a maximum target length of $88$. Our results are slightly higher than the former and significantly higher than the latter. It is worth noting that both studies leverage a number of samples at least an order of magnitude higher than ours. In addition, the function lengths analyzed with our approach are greater than those used by either work. For the sake of clarity, we note that these works adopt a complementary version of ED, less widespread for NMT tasks (liguori2023evaluates, ). We provide further details about these works in Section 6. Focusing on the other metrics obtained from fairseq, ROUGE-L shows that, in 77% of cases, the generated code contains word sequences similar to those of the reference. We observe a 4-gram overlap of 58% (BLEU-4) and a token order adherence of 64% (METEOR).
For the sake of completeness, we also evaluated the performance of a traditional decompiler, Ghidra (Ghidra, ). Specifically, we randomly sampled 1,000 programs from our decompiled test set. Ghidra relies on control flow graphs (CFGs), data flow graphs (DFGs), and other intermediate representations (IRs) to perform decompilation. Ghidra results, with a 23% ED, indicate significantly lower performance compared to studies utilizing neural network-based models, which suggests that traditional graph-based methods may struggle to capture the complexities of the source code as effectively as neural approaches. However, unlike a traditional decompiler, an NMT model always generates an answer despite having a low probability of approaching the correct one. In addition, one of the disadvantages of using a neural network is that there is less control over the translation, i.e., in case of errors it is very complex to determine which factor led the model to a specific decision over others. An interesting finding is that pre-trained models, i.e., CodeT5+ and CodeBERT, perform considerably worse than the model trained from scratch, i.e., fairseq. Above all, CodeBERT presents a discrepancy of around 30% on BLEU-4 and ROUGE compared to fairseq. In most of the metrics CodeT5+ scores about 10% lower than fairseq, reaching in some cases intermediate values between fairseq and CodeBERT, as in the case of the ED (49%). As noted by Tufano et al. (tufano2023automating, ), we can associate the negative impact of pre-training with the size of the training dataset. Pre-training helps boost performance only if the amount of fine-tuning data available is small. We deem our dataset with about 75,000 samples large enough to overwrite the parameters tuned in the pre-training phase. According to these results, we adopt fairseq as the neural decompiler for further analysis.
Table 5. Qualitative analysis on decompilation.
*: the results on the multi-class classification only concern Flawfinder-compatible CWEs. °: the results on the multi-class classification concern different CWEs selected by Schaad et al. (schaad2023deep, ).
Binary Classification. The first analysis concerns the classification of the target code as vulnerable/non-vulnerable. We submitted the C/C++ code translated by the neural decompilation model to the classification process, adopting the same data pre-processing for each model. We selected accuracy and F1-Score as the metrics to evaluate the performance of the models.
Table LABEL:tab:Classifications shows the models’ performance for the binary classification on the decompiled C/C++ code. CodeGPT is the best-performing model with an accuracy and F1-Score of 95%. The performance is high for all models, with both metrics around 90%. The highest gap lies between SRNN and CodeGPT, exhibiting a discrepancy of about 7% in terms of F1-Score. Table LABEL:tab:Classifications also reports the performance of a state-of-the-art work on vulnerability prediction in binary files (schaad2023deep, ), which achieves an 88% F1-Score through the LLVM IR and the SRNN model, further described in Table 7 and Section 6.
Furthermore, we tested a static analysis tool, Flawfinder (flawfinder, ), on the same test set. This tool scans C/C++ source code and flags potential security holes. Flawfinder exhibits metrics that are close to 50% on average, which is equivalent to a random model. Therefore, the results support the implemented design choice of using DL models for vulnerability prediction instead of a static analysis tool. Finally, we leveraged Shap (NIPS2017_7062, ), a tool for explaining the output of our vulnerability prediction models, to better understand their behavior and which features are the most influential for the learning process, i.e., the tokens most exploited for classification. This is a helpful analysis to avoid falling into the so-called Clever Hans effect, which occurs when the model produces correct predictions based on ”wrong” features (kauffmann2020clever, ). The results revealed tokens in C/C++ code such as: if, new, strcat, funct, static, malloc, source, and printline. These results allow us to state that the model focuses on tokens that are relevant to implementing countermeasures to vulnerabilities, e.g., if constructs to check specific conditions (e.g., underflow or overflow), without taking decisions based on ”easier” alternatives, for instance, the length of the program under analysis (excluded by padding mechanism).
Table 7. Comparison of vulnerability prediction works on binary files.
Work Dataset Classes of Standardized Intermediate Vulnerability Binary Multi-class
Vulnerabilities Taxonomy Representation Predictor Classification Classification
Accuracy Accuracy
Pewny et al. (pewny2015cross, ) OpenSSL 5 IR+CFG Greedy 61%
DiscovRE (eschweiler2016discovre, ) Github 2 CFG k-NN 93%
Dahl et al. (dahl2020stack, ) Open-source Code 1 ASM Tokens RNN 99%
VDISCOVER (grieco2016toward, ) Debian 4 Call Graph Random Forest 55%
Zheng et al. (zheng2019recurrent, ) SARD 4 CWE LLVM IR RNN 90%
Schaad et al. (schaad2023deep, ) Juliet 23 CWE LLVM IR RNN 88% 77%
Proposed solution Juliet 20 CWE Decompiled C/C++ Transformers 95% 83%
Multi-class Classification. This task aims to identify the CWE associated with a vulnerable function. Specifically, the number of classes (CWEs) for this classification task is 20, plus an additional class associated with non-vulnerable (good) functions. The complexity of this task is related to the similarity between test cases of different classes. Therefore, a deep inspection of the test cases is necessary to determine the actual CWE to which they belong. We decided not to select classes based only on the highest number of occurrences in the dataset. Instead, we manually selected classes based on the dissimilarity between different CWEs, to reduce the confusion degree for the model towards similar CWEs. Consequently, to the best of our knowledge, we evaluated the difference between CWEs from a logical point of view. For instance, we selected one CWE between 190 Integer Overflow and 191 Integer Underflow because of their semantic similarity to prevent the situation in which the model hardly distinguished one from the other. Moreover, we discarded CWEs 546 Suspicious Comment and 615 Inclusion of Sensitive Information in Source Code Comments since all comments were removed during the pre-processing stage, described in Section 3.3. It is worth noting that each class needs a significant number of test cases so that the model can learn from as many examples as possible. The requirement on the number of tests is that each class has at least $300$ samples (at least $150$ bad and $150$ good examples). We prioritized the CWEs that belonged to the Top 25 CWEs published yearly by MITRE among the most represented ones, i.e., those with a reasonable amount of samples.
Table LABEL:tab:Classifications shows the accuracy and F1-Score of the models for the multi-class classification task. The models’ performance decreased from the binary classification case, with CodeBERT performing the best with an accuracy of 83% and F1-Score of 82%. The lowest performance is again achieved by SRNN with 70% F1-Score, showing a discrepancy from CodeBERT of about 12%. The performance gap between our approach, utilizing CodeBERT, and the method proposed by Schaad et al. (schaad2023deep, ) is approximately 6% in terms of accuracy and 3% in the F1-score. However, it is important to note that the 20 classes analyzed in our study differ both in type and quantity from their 23 classes (schaad2023deep, ), as shown in Table 7.
Flawfinder (flawfinder, ) is compatible with only a subset of CWEs (CWEflawfinder, ). Consequently, we conducted an experiment excluding all non-compatible CWEs from our dataset, resulting in approximately 500 programs. The performance in this test was notably worse compared to our earlier results on the binary classification task, highlighting the potential advantage that deep learning techniques could offer in this context.
Table 7 highlights the key improvements of the proposed approach in its entirety compared to similar works. Notably, in addition to achieving higher performance, our method uniquely leverages the transformer architecture for both tasks and includes source code as an intermediate representation. Therefore, we can conclude that relying on source code as IR, instead of disassembled code, resulted in a 7% increase in bi-class analysis performance and a 6% improvement in multi-class.
Qualitative analysis. We performed a qualitative analysis to provide deeper insight into the vulnerability prediction evaluation. Specifically, we refer to the CodeBERT model, i.e., the best-performing model in the multi-class prediction. Table 4.4 shows two vulnerable examples of multi-class classification. The first is an example of misprediction: the program associated with CWE 369 Divide By Zero was incorrectly associated with CWE 197 Numeric Truncation. The original program presented the instruction int result = (int)(100.0/data), i.e., the core of the vulnerability, completely omitted in the decompiled version. Despite a high ED metric, this error propagated into the prediction task, causing the model to fail.
The second example displays a correct prediction for CWE 190 Integer Overflow. In this case, the model successfully identifies the vulnerability type, leveraging the high quality of the decompiled program (ED equal to 88%).
4.4. Impact of decompilation quality on vulnerability prediction
The proposed solution leverages the output of the neural decompilation as the input for the vulnerability prediction. Consequently, it is both valuable and insightful to investigate how the quality of the decompiled code influences the performance of the subsequent classification model.
Table 8. Qualitative analysis on multi-class classification.
Example 1 Example 2
Decompiled C/C++ void funct 0 () { int data ; myUnion ; data = - 1 ; fscanf ( stdin , " ", & data ); myUnion . unionFirst = data ; { int data = myUnion . unionSecond ; { short shortData = ( short ) data ; printShortLine ( shortData ); } } } static int staticFive = 5 ; void funct 0 () { unsigned int data ; data = 0 ; if ( staticFive == 5 ) { data = (unsigned int) RAND 3 2 (); } if ( staticFive == 5 ) { { unsigned int result = data - 1 ; printUnsignedLine ( result ); } } }
ED 70% 88%
CWE observed CWE 369: Divide by Zero CWE 190: Integer Overflow
CWE predicted CWE 197: Numeric Truncation (misprediction) CWE 190: Integer Overflow (correct prediction)
Figure 3. Bar Chart showing the impact of neural decompilation quality (ED-Mean) on vulnerability prediction (F1-Score) for each classified CWE.
Figure 3 illustrates the comparison between the F1-Score and ED-Mean (Edit Distance Mean) across the classified CWE categories. It is worth noting that CWEs as 762, 36, and 23 present high values for both ED-Mean F1-Score. In contrast, CWE 197 presents extremely low values for both F1-Score and ED-Mean. To investigate the relationship between the models’ performance on these tasks we analyze the correlation between the two phenomena. We adopted Pearson’s correlation coefficient, denoted as $r$, with $x$ as the F1-Score, $y$ as the ED-Mean, and $n$ as the number of $(x,y)$ pairs (i.e., the number of CWE entries in this case). The computed correlation coefficient between the F1-Score and the ED-Mean is:
$r=\frac{n\left(\sum xy\right)-\left(\sum x\right)\left(\sum y\right)}{\sqrt{\left[n\sum x^{2}-\left(\sum x\right)^{2}\right]\left[n\sum y^{2}-\left(\sum y\right)^{2}\right]}}=\textbf{0.65}$
The result indicates a moderate, positive correlation. This suggests that as the ED-Mean improves, the average F1-Score value tends to increase correspondingly, reflecting a relationship where higher classification performance is associated with smaller error distances.
On the reverse, as demonstrated in the first example in Table 4.4, it is still plausible that a high ED score may lead to a low F1-Score. Even though the translation may present a high quality standard, two kinds of issues may still arise: i) the decompiled code may be logically flawed, ii) the vulnerability prediction process may have failed to identify the underlying vulnerability.
5. Lessons Learned
The results of the experimental study provided us with several lessons learned about using neural decompilation in the cybersecurity landscape.
Applicability of neural decompilation to cybersecurity. The idea of leveraging neural decompilation for cybersecurity-related tasks is promising. Specifically, given the quality of the translated code, neural decompilation showed great potential for vulnerability prediction, with the possibility to unlock several novel applications, e.g., binary malware recognition. We analyzed three state-of-the-art models, i.e., fairseq, CodeBERT, and CodeT5+, with fairseq emerging as the best-performing model, achieving quality translations comparable to other studies in code translation and providing a representative basis for our experiments. Future work includes integrating additional models for neural decompilation, with a deeper inspection of pre-trained NMT models.
Use of DL for vulnerability prediction. The results of the vulnerability prediction task revealed that DL models were indeed appropriate for vulnerability detection and identification, with transformer-based models emerging as the best performing. The selected models could accurately predict whether a C/C++ function was vulnerable and the related vulnerability in the MITRE CWE taxonomy.
6. Related Work
Neural decompilation is a recent and innovative research trend in binary analysis. Hosseini et al. (hosseini2022beyond, ) introduced Beyond The C, a neural decompiler based on the fairseq model. This solution is retargetable for several high-level programming languages such as C, OCaml, Go, and Fortran. The results of their study are promising for all languages, scoring the best result in terms of ED ($54\%$) on the decompilation of C/C++ code (training on $2,000,000$ samples). Katz et al. (katz2018using, ) proposed an approach for neural decompilation using RNNs. The solution was trained and evaluated on snippets of binary machine code compiled from C source code. This solution achieved an ED equal to $30\%$, leveraging about $700,000$ samples.
The neural decompilation phase of our experimental study builds on top of the solution proposed by Hosseini et al. but differently, we selected a different set of programs and performed a comparative analysis between state-of-the-art NMT models. In addition, rather than focusing on neural decompilation only, our study is focused on applying this novel solution to the cybersecurity landscape, in particular vulnerability prediction on binary code.
Zheng et al. (zheng2019recurrent, ) devised an approach to identify vulnerabilities in binary files using deep learning (DL) models, in particular Recurrent Neural Networks (RNN). To this end, this work leverages the Juliet C/C++ dataset (Juliet, ), described in Section 2.2. Then, the binary code is lifted to LLVM (LLVM, ) intermediate representation (IR), constituting the input for the RNNs. This solution can identify 4 CWEs with 90% accuracy. Schaad et al. (schaad2023deep, ) extended the previous approach (zheng2019recurrent, ), leveraging several DL models, i.e., Recurrent Neural Networks (RNN), Long-Short Term Memory (LSTM), and Gated Recurrent Unit (GRU). Their solution can identify if a binary file is vulnerable with 88% accuracy, and classify 23 CWEs out of the 118 described in the Juliet dataset with 77% accuracy. Pewny et al. (pewny2015cross, ) designed a cross-architecture solution to identify bug signatures in binary executables: the binaries are lifted to IR to unify the different assembly languages and then converted to control flow graphs (CFG) to compare basic blocks to the known signatures. This solution can identify 5 technology-specific vulnerabilities across three architectures (ARM, MIPS, x86), including the well-known Heartbleed vulnerability for OpenSSL (HeartbleedCVE, ; Heartbleed, ). DiscovRE (eschweiler2016discovre, ) is a prototype for identifying binary code bugs. It can identify two vulnerabilities, i.e., Heartbleed and POODLE for OpenSSL and SSLv3 respectively, and supports four architectures, namely x86, x64, ARM, and MIPS. DiscovRE measures structural similarities between the target code and the vulnerable prototype through custom numeric features and metrics to identify the vulnerabilities. Dahl et al. (dahl2020stack, ) developed a solution to identify stack-based buffer overflow vulnerabilities in binary files, along with a custom dataset for this task. First, the binary code is disassembled and then compressed to discard the redundant information. Again, the vulnerability detection model is an LSTM, which achieves 90% accuracy in detecting stack-based buffer overflows. VDISCOVER (grieco2016toward, ) is a ML-based tool for memory corruption detection. It comes with a custom dataset, VDISCOVERY, built from 1,039 programs of the Debian bug tracker. VDISCOVER can identify 4 types of memory corruptions, namely heap memory, stack memory, corruption of argument in a function, and corruption of a pointer to a function, with 55% accuracy. Table 7 summarizes the comparison between the related work and our solution. Our work differs from the previous ones, in terms of binary code analysis and vulnerability prediction. On the one hand, neural decompilation does not need any formal representation of the binary code (e.g., control flows, intermediate representation) since the code is treated as plain text for the decompilation task. The use of NMT for this task reduces the complexity of the decompilation process, requires less knowledge about the source and target languages, and makes the solution language independent, i.e., it is potentially applicable to a different programming language or an alternative target architecture with the only cost of re-training the model. Indeed, we can observe that only two works, namely DiscovRE (eschweiler2016discovre, ) and the work from Pewny et al. (pewny2015cross, ), are retargetable for the architecture. On the other hand, contrary to other works that focus on technology-specific or single vulnerabilities, we focus on identifying several vulnerabilities described in standard taxonomies, namely the CWE provided by MITRE, achieving higher accuracy than two related works (pewny2015cross, ; grieco2016toward, ) and also surpasses the only work that encompasses a comparable number of vulnerabilities (schaad2023deep, ). Even though the other works achieve a slightly better performance, they only analyze a significantly smaller number of technology-specific vulnerabilities (zheng2019recurrent, ; eschweiler2016discovre, ; dahl2020stack, ). Using a standardized vulnerability taxonomy is a key aspect in evaluating these works: technology-specific vulnerabilities are harder to generalize and typically do not apply to different systems, while standardized vulnerabilities, e.g., CWEs, represent global definitions that span several kinds of software systems and products.
7. Threats to Validity
Dataset. The chosen dataset, Juliet C/C++ 1.3, is a synthetic dataset, i.e., novel data created from real data with the same statistical properties. Using a synthetic dataset provides a large amount of samples, accurately annotated with ground truth. However, the synthetic data do not perfectly align with actual data regarding the size and complexity of real code. Therefore, the vulnerability detection accuracy on synthetic data may be higher than in real data. This may hinder the external validity of the study. Nevertheless, the synthetic data still allow us to analyze the decompilation task and the prediction of vulnerabilities in the translated code, which represents a good starting point for this research direction. Moreover, it is challenging to find datasets in the literature that satisfy the requirements for vulnerability prediction in binary code, as illustrated in Section 2.2. Specifically, Juliet represents a fair trade-off between the requirements for the dataset: i) a dataset composed of compilable C/C++ functions and ii) labeled with the vulnerabilities from a standardized taxonomy.
Models selection. The selection of transformer-based and DL models for the neural decompilation and vulnerability prediction tasks may impact the study’s external validity. On the one hand, we selected three state-of-the-art models, namely fairseq, CodeBERT, and CodeT5+, for neural decompilation performing a comparative study to assess their capabilities for this task. On the other hand, we leveraged CodeBERT, CodeT5+, CodeGPT, and traditional DL models for vulnerability prediction since they represent the most popular solution for this task (zheng2019recurrent, ; schaad2023deep, ; dahl2020stack, ; chen2023diversevul, ). We tried to include, to the best of our knowledge, models with different architectures and capabilities, representative of the state-of-the-art.
Metrics. To evaluate the performance of neural decompilation, we selected output similarity metrics such as BLEU, METEOR, ED, and ROUGE. This choice may threaten the construct validity since such metrics may not capture the correctness of the decompiled code. Nevertheless, output similarity metrics represent the most popular practice for evaluating NMT models, both for code-to-code (as in similar works on neural decompilation (hosseini2022beyond, ; katz2018using, )) and NL-to-code translation tasks (liguori2023evaluates, ). Specifically, ED represents the prevalent metric to estimate the performance of code-to-code translation.
8. Conclusion
In this work, we presented an experimental study on vulnerability prediction in binary files. In particular, we leveraged neural decompilation to translate binary code into source code, treating both programs as simple text. Our goal was to demonstrate that a more effective solution for detecting vulnerabilities in binary files is to use neurally decompiled source code, rather than relying on complex toolchains to convert binary code into an intermediate representation or to extract control flows for decompilation. Our experimental analysis revealed that neural decompilation is fit for this task, achieving great results comparable with state-of-the-art works, and that transformer-based and DL models can accurately predict whether a binary file is vulnerable and identify the type of vulnerabilities, mapped to the MITRE CWE taxonomy, using the neurally decompiled code.